Merck & Co., Inc. is a world leader in the discovery, development, manufacture and marketing of a

broad range of human and animal health products. The company, which has 70,000 employees, spends over USD 2

billion every year on the research and development of new drugs. As of the end of 2, its 2.2 billion shares are valued

in the stock market for a total of USD 132 billion. Given the following data for Merck, calculate the equity ratios for

2003 and 2002. Then comment on the results.

2003 2002

Stockholders' equity USD 14,832,400,000 USD 13,241,600,000

Total equities USD 39,910,400,000


1
Expert's answer
2021-07-12T15:53:07-0400

Solution:

The formula for equity ratio is as follows:


Equity ratio = "\frac{Total \;shareholders\; equity}{Total\; assets}"


Total equity is the same as total assets


Equity ratio year 2002 = "\frac{13,241,600,000}{35,634,900,000} = 0.37"

 

Equity ratio year 2003 = "\frac{14,832,400,000}{39,910,400,000} = 0.37"

 

From the above calculations of equity ratios for the year 2002, and 2003, the equity ratios derived are the same and are below 0.5. This means that Merck & Co., Inc. is a leveraged company. That is, the company is funded by more debts than equity since it holds more debts.